Draw the cumulative frequency curve for the data given below
Marks: 0-10 10-20 20-30 30-40
No of students: 4 8 11 15
40-50 50-60 60-70
12 6 3
1
Expert's answer
2021-07-26T08:03:12-0400

The total number of students: "N=4+8+11+15+12+6+3=59"

Therefore, the corresponding partial frequencies (xi) are 4/59, 8/59, 11/59, 15/59, 12/59, 6/59, 3/59

The accumulated relative frequencies can be found by summarizing "S_i=x_1+\dots +x_i" or by the formula "S_{i}=S_{i-1}+x_i".


"S_1=4\/59", "S_2=(4+8)\/59=12\/59", "S_3=(12+11)\/59=23\/59", "S_4=(23+15)\/59=38\/59", "S_5=(38+12)\/59=50\/59", "S_6=(50+6)\/69=56\/59",

"S_7=(56+3)\/59=1"


Put the obtained midpoints of the intervals, relative frequencies and accumulated relative frequencies into the table:



Draw the cumulative frequency curve: